OpenStudy (lovelyanna):

When rolling a number cube, what is the probability of rolling a number less than 3? A. 1/6 B. 1/3 C. 1/2 D. 2/3 I think B

OpenStudy (cwrw238):

the result will be 1 or 2 probabiilty of a 1 is 1/6 as there are 6 numbers on the dice and same fot the 2 , 1/6 for 1 OR 2 we add the probabilities = 1/6 + 1/6

OpenStudy (anonymous):

6 number there is 2 under 3

OpenStudy (cwrw238):

B is correct

OpenStudy (anonymous):

2/6

OpenStudy (anonymous):

1/3

OpenStudy (lovelyanna):

Thought so, thank you both!

OpenStudy (cwrw238):

yea another way to look at is there are 6 ways to get 1 - 6 there are 2 ways to get 1 or 2 so probabiliity of 1 or 2 = 2/6 = 1/3

OpenStudy (lovelyanna):

Mmm, I see!